OpenStudy (anonymous):

A canoeist paddles 21 km downstream in 3h. The return trip takes 1.2 h longer. What is the rate of the current?

OpenStudy (anonymous):

You have to know two things: The area of the flow and the speed of the flow. To get the area you have to measure the depth and width of the river at a particular cross-section perpendicular to the flow direction.

OpenStudy (phi):

Do you know rate * times = distance or velocity * time = distance or speed * time= distance (All mean the same thing) Let's use rate. The rate of the paddler is P (for paddler) in km/hr (kph) in still water Call the rate of the current C (C for current) if she is going with the current we add in how fast the current is: downstream the rate is P+C going upstream the paddler goes slower because the current subtracts from her rate, upstream the rate is P-C now use rate * time = distance can you write down the equation when going downstream ?

OpenStudy (anonymous):

21km/3h?

OpenStudy (phi):

yes, but that is not an equation. I meant like this: DOWNSTREAM: rate * time = distance replace rate with (P+C) (that is the rate going downstream, though we don't know P or C) replace time with 3 hours replace distance with 21 km we get (P+C) * 3 hrs = 21 km that I is what I mean by "write the equation" can you write down the equation for going upstream?

OpenStudy (anonymous):

Oh sorry , upstream = (P-C)4.2hrs=21km

OpenStudy (phi):

yes, looks good. now we use algebra to solve for C to answer the question What is the rate of the current? so we have 3(P+C) =21 4.2(P-C)= 21 I would simplify both equations. In the first, divide both sides by 3 (like you posted up above) . In the 2nd, divide both sides by 4.2 what are the equations now?

OpenStudy (anonymous):

(P+C)=7km/h (P-C)=5km/h

OpenStudy (phi):

to find C, I would subtract: top equation minus bottom equation. Can you finish ?

OpenStudy (anonymous):

2P=12km/h P=6km/hr 6km/hr+C=7km/hr C=1km/h

OpenStudy (phi):

Yes if you added the 2 equations you would get (P+C)=7km/h (P-C)=5km/h 2P +C -C = 12 2P= 12 P= 6 and then you can find C you could also subtract equations like this (P+C)=7km/h -(P-C)=-5km/h (put a minus sign in front of both sides) P+C-(P-C) = 7 - 5 P+C-P +C = 2 2C= 2 C=1 km/hr

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

yw